# To See a World in a Spark of Neuron: Disentangling Multi-task Interference for Training-free Model Merging

### v2

@@ Abstract
Fine-tuning pre-trained models on targeted datasets enhances task-specific performance but often comes at the expense of generalization. Model merging techniques, which integrate multiple fine-tuned models into a single multi-task model through task arithmetic, offer a promising solution. However, task interference remains a fundamental challenge, leading to performance degradation and suboptimal merged models. Existing approaches largely overlook the fundamental roles of neurons, their connectivity, and activation, resulting in a merging process and a merged model that does not consider how neurons relay and process information. In this work, we present the first study that relies on neuronal mechanisms for model merging. We decompose task-specific representations into two complementary neuronal subspaces that regulate neuron sensitivity and input adaptability. Leveraging this decomposition, we introduce NeuroMerging, a novel merging framework developed to mitigate task interference within neuronal subspaces, enabling training-free model fusion across diverse tasks. Through extensive experiments, we demonstrate that NeuroMerging achieves superior performance compared to existing methods on multi-task benchmarks across both natural language and vision domains. Our findings highlight the importance of aligning neuronal mechanisms in model merging, offering new insights into mitigating task interference and improving knowledge fusion. Code will be released upon acceptance.
To See a World in a Spark of Neuron: Disentangling Multi-task Interference for Training-free Model Merging
Zitao Fang1, Guodong Du2, Shuyang Yu1, Yifei Guo1, Yiwei Zhang1, Yiyao Cao1, Jing Li2, Ho-Kin Tang2*, Sim Kuan Goh1* 1Xiamen University Malaysia, 2Harbin Institute of Technology, Shenzhen, China ait2209071@xmu.edu.my, denghaojian@hit.edu.cn, simkuangoh@gmail.com
@@ 1 Introduction
Model Merging began with weight interpolation Wortsman et al. (2022) to combine the strengths of different models in weight space by balancing competition and cooperation within shared representation Du et al. (2024); Ilharco et al. (2023); Ortiz-Jimenez et al. (2023). In NLP, methods such as merging task-specific language models have been explored to build and update foundation models with multi-task capabilities Raia (2021); Wan et al. (2024a); Akiba et al. (2025); Wan et al. (2024b). Similarly, in CV, approaches like merging Vision Transformers (ViTs) trained on different tasks or domains have been investigated to create unified models capable of handling diverse visual tasks Kim et al. (2021); Bao et al. (2022); Wang et al. (2024). In the multi-modal space, model merging has been applied to integrate models from different modalities, such as text and images, enhancing tasks like audio-visual question answering and image captioning Sung et al. (2023); Sundar et al. (2024); Dziadzio et al. (2025). These advancements underscore model merging as a promising avenue for future research.
Existing methods for model merging primarily operate at three granularities: model-level, layer-level, or parameter-level Ilharco et al. (2023); Yang et al. (2024b); Du et al. (2024), while overlooking the fundamental roles of neurons, their activation and connectivity Suhaimi et al. (2022); Stelzer et al. (2021), which underpins the learning process all neural networks from Perception Rosenblatt (1958) to LLMs Touvron et al. (2023). Appendix Table 6 provided a summary of existing methods across different granularities. In Figure 2 and Appendix Table 6.2, we analyze on how changes in model weights along two complementary neuronal subspaces, leads to distinct impacts on task performance, by removing one and retaining the other. Notably, one subspace preserves most of the task-specific capabilities. This observation motivates us to explore model merging at the neuronal level, which could have important implications for mitigating task interference and could yield more robust merged models.
@@
Figure 3: Illustration of our proposed framework. Neuronal Task Vector: Neuronal task vectors $\tau$ for the $k^{th}$ neuron are defined as the difference between the fine-tuned and pre-trained neuron for each task. Decomposition: The pre-trained $k^{th}$ neuron is decomposed into its parallel and orthogonal complementary subspaces, followed by the projection of neuronal task vectors onto these subspaces. NeuroMerging: Our proposed NeuroMerging operates within these complementary subspaces for neuronal model merging.
@@ 2 Related Work
Multi-task learning (MTL) Fifty et al. (2021) leverages transferable knowledge to handle multiple related tasks simultaneously. Existing MTL approaches primarily rely on architectural design or optimization strategies. Architectural-based methods, such as Mixture of Experts (MoE) Shazeer et al. (2017), introduce specialized subnetworks that dynamically route inputs to task-specific experts, effectively reducing interference. However, these methods require modifying the pretrained model structure, increasing computational complexity, and limiting scalability Liu et al. (2019); Shen et al. (2024). Optimization-based approaches, on the other hand, focus on balancing task gradients or loss functions to mitigate task conflicts during training Bai et al. (2023); Cipolla et al. (2018). While these methods improve convergence, they still depend on task-specific training data, which may be impractical in real-world applications due to privacy concerns or data scarcity Liang et al. (2022). In contrast, model merging offers an alternative paradigm by integrating knowledge from multiple fine-tuned models into a single unified model without requiring additional training data or architectural modifications Wortsman et al. (2022); Ilharco et al. (2023). Notwithstanding the promising findings, a key challenge in model merging is task conflict Yadav et al. (2023); Du et al. (2024), where different tasks compete for model capacity, potentially leading to suboptimal performance.
@@ 3 Methodology
In this section, we first formalize the concept of neuronal task vector for model merging and then decompose neuronal task vectors into two complementary neuronal subspaces. Subsequently, we introduce our framework, NeuroMerging, showed in Algorithm 1 in Appendix A, which performs merging in the neuronal subspaces.
@@ 3.4 NeuroMerging
where $\psi_{\parallel}(\cdot)$ denotes the merging function for $\tau^{k}_{\parallel,t}\in\mathbb{R}$, which can be a commonly used weighted average, TIES’s disjoint merge Yadav et al. (2023), and others. $\psi_{\parallel}(\cdot)$ is also applied to non-neuronal parameters such as bias and pre-norm. For $\tau^{k}_{\perp,t}\in\mathbb{R}^{T}$, we first find dominant orthogonal subspaces within $\mathcal{O}$ using singular value decomposition (SVD) with rank being the number of tasks interact within the same neuron, and then project $\tau^{k}_{\perp,t}$ along each dimension of the SVD subspace, before applying $\psi_{\parallel}(\cdot)$ to them.
$\lambda_{1}$ and $\lambda_{2}$ are scaling parameters. When validation data is available $\lambda_{1}$ and $\lambda_{2}$ could be tuned. However, when validation is unavailable, we propose setting $\lambda_{1}=0$ (as the corresponding subspace is observed to have little impact in Section 6.3) and estimate $\lambda_{2}$ based on the impact of top $r\%$ mentioned in Section 3.1. Specifically, $\lambda_{2}$ is estimated as $\lambda_{2}=\frac{1}{1-\sigma}$, where $\sigma=max(\sigma_{1},...,\sigma_{t},...,\sigma_{T})$ and $\sigma_{t}$ is the ratio of the $L_{1}-Norm$ of the zeroed-out elements in the task vector in Section 3.1 to the $L_{1}-Norm$ of the original task vector: $\sigma_{t}=\frac{\|\mathbf{\tau}_{t}^{masked}\|_{1}}{\|\mathbf{\tau}_{t}\|_{1}}$. Detailed discussion on the parameters is provided in Section 6.3.
Subsequently, we reconstruct the task vector $\overline{\tau}$ with the all the merged neuronal task vectors $\overline{\tau}_{k}$. Finally, we obtained the final merged model $\overline{\theta}=\theta_{0}+\overline{\tau}_{rescaled}$. Additional discussion is provided in Appendix B.
@@ 4 Experimental Setup
Baseline Methods. Our baselines comprise two main categories: (1) non-model merging approaches, which include individually fine-tuned models and a multitask model trained jointly on the combined dataset serving as our theoretical upper bound, and (2) various advanced model merging techniques, including Simple Averaging (Wortsman et al., 2022), Fisher Merging (Matena and Raffel, 2022), RegMean (Jin et al., 2023), Task Arithmetic (Ilharco et al., 2023), TIES-Merging (Yadav et al., 2023), AdaMerging (Yang et al., 2024b), PCB-Merging (Du et al., 2024), and our proposed NeuroMerging method. Notably, we use task-wise AdaMerging for fair comparison with existing baselines and our method. We report average accuracy across all tasks’ test sets as our primary evaluation metric.
@@
Validation Set Availability. Previous works exhibit varying dependencies on a validation set. Fisher Merging inherently requires a validation set to compute the Fisher matrix. Other approaches may optionally utilize validation data for hyperparameter tuning, while RegMean leverages training data to compute and store inner product matrices for model merging. However, since these matrices match the dimensions of the original model, they introduce substantial storage and computational overhead, limiting scalability to larger models and more extensive merging tasks.
@@
Hyperparameters. In the absence of an additional validation set, we set $\lambda=1$ as the default value for all task-vector-based methods. For TIES-Merging and PCB-Merging, which require a masking ratio, we follow the settings of Yadav et al. (2023) and Du et al. (2024), applying $r=0.2$ as the default value across all experiments. For NeuroMerging, we set a default masking ratio of $r=0.15$, with $\lambda_{1}$ fixed at 0, while $\lambda_{2}$ is automatically adjusted according to the methodology described in Section 3.4.
When validation is allowed, we configure the non-diagonal multiplier $\alpha$ in RegMean to $0.9$, except for the T5-base model, where it is set to $0.1$. For Task Arithmetic, we perform a grid search over $\lambda$ ranging from $0.2$ to $1.5$ with a step size of $0.1$. For TIES-Merging, PCB-Merging, and NeuroMerging, we search for the optimal masking ratio r in the range $[0.05,0.2]$ with a step size of $0.05$, and $\lambda$ ($\lambda_{2}$ for NeuroMerging) from $0.8$ to $5.0$ with a step size of $0.1$.
@@ 5 Results
5.1 Merging NLP Models
Following the experimental settings from Yadav et al. (2023), we use the T5-base and T5-large models (Raffel et al., 2020), which are encoder-decoder transformers (Vaswani et al., 2017) pretrained via masked language modeling on a large text corpus, and fine-tune them independently on seven tasks: Khot et al.’s (2019) QASC, Yang et al.’s (2015) WikiQA, and Tafjord et al.’s (2019) QuaRTz for Question Answering; Zhang et al.’s (2019) PAWS for Paraphrase Identification; Sharma et al.’s (2018) Story Cloze for Sentence Completion; and Sakaguchi et al.’s (2021) Winogrande together with Levesque et al.’s (2012) WSC for Coreference Resolution. Table 2 and 12 demonstrated that our approach has superior performance over state-of-the-art methods, achieving improvements of $0.6\%$ and $1.2\%$ for T5-base and T5-large, respectively. Moreover, NeuroMerging without validation showed even more substantial gains, surpassing previous methods by $1.4\%$ for T5-base and $0.7\%$ for T5-large. For comprehensive results across all tasks and model variants, see Appendix C.3.
5.2 Out-of-Domain Generalization
Building upon the experimental setup of Yadav et al. (2023), we also investigated how merged models between tasks enhance generalization in different domains. Following the approach used in prior NLP models, we merged models on seven in-domain datasets and evaluated their performance on six held-out datasets from the T0 mixture (Sanh et al., 2022) to assess cross-task generalization. These datasets encompass diverse tasks, covering three Question Answering datasets: Huang et al.’s (2019) Cosmos QA, Sap et al.’s (2019) Social IQA, and Rogers et al.’s (2020) QuAIL; one Word Sense Disambiguation dataset: Pilehvar and Camacho-Collados’s (2019) WiC; and two Sentence Completion datasets: Gordon et al.’s (2012) COPA and Zellers et al.’s (2019) H-SWAG. As shown in Figure 4, NeuroMerging outperformed the strongest baseline by $0.7\%$ and $0.5\%$ for T5-base and T5-large models, respectively, showcasing enhanced out-of-domain generalization capabilities. For more comprehensive results, please refer to Appendix C.3 Table 15 and 16.
@@ 5.3 Merging LLMs
We follow the experimental setup of Du et al. (2024) and extend our NeuroMerging to larger LLMs. Specifically, we merged three domain-specialized Llama-2-7b models (Touvron et al., 2023), and each is fine-tuned for distinct capabilities: Chinese language understanding, mathematical reasoning (Yu et al., 2024b), and code generation (Roziere et al., 2023). To rigorously evaluate the performance of each specialized model, we employed established benchmarks tailored to their respective domains: Li et al.’s (2024) CMMLU for assessing Chinese language proficiency, Cobbe et al.’s (2021) GSM8K for mathematical capabilities, and Chen et al.’s (2021) HumanEval for code generation competency. As demonstrated in Table 3, our NeuroMerging approach exhibited substantial performance improvements, surpassing the strongest baseline by $0.9\%$. Notably, our method even outperformed the PCB-Merging utilizing Evolution Strategies (ES) optimization algorithm by $0.7\%$, underscoring the effectiveness of our proposed methodology.

### v4

@@ Abstract
Fine-tuning pre-trained models on targeted datasets enhances task-specific performance but often comes at the expense of generalization. Model merging techniques, which integrate multiple fine-tuned models into a single multi-task model through task arithmetic, offer a promising solution. However, task interference remains a fundamental challenge, leading to performance degradation and suboptimal merged models. Existing approaches largely overlooked the fundamental roles of neurons, their connectivity, and activation, resulting in a merging process and a merged model that does not consider how neurons relay and process information. In this work, we present the first study that relies on neuronal mechanisms for model merging. Specifically, we decomposed task-specific representations into two complementary neuronal subspaces that regulate input sensitivity and task adaptability. Leveraging this decomposition, we introduced NeuroMerging, a novel merging framework developed to mitigate task interference within neuronal subspaces, enabling training-free model fusion across diverse tasks. Through extensive experiments, we demonstrated that NeuroMerging achieved superior performance compared to existing methods on multi-task benchmarks across both natural language and vision domains. Our findings highlighted the importance of aligning neuronal mechanisms in model merging, offering new insights into mitigating task interference and improving knowledge fusion. Our project is available at https://ZzzitaoFang.github.io/projects/NeuroMerging/.
To See a World in a Spark of Neuron: Disentangling
Multi-task Interference for Training-free Model Merging
Zitao Fang1  Guodong Du2,*  Shuyang Yu3,†  Yifei Guo4,† Yiwei Zhang1Yiyao Cao1Jing Li5Ho-Kin Tang5Sim Kuan Goh1,* 1Xiamen University Malaysia  2The Hong Kong Polytechnic University 3Columbia University  4Duke University  5Harbin Institute of Technology (Shenzhen) ait2209071@xmu.edu.my, duguodong7@gmail.com, simkuangoh@gmail.com
11footnotetext: Corresponding authors.22footnotetext: Work conducted while at Xiamen University Malaysia.
@@ 1 Introduction
Method Scale Granularity Level
Fisher Merging[NeurIPS22] Fisher Matrix Parameter
RegMean[ICLR23] Inner Product Matrix Parameter
Task Arithmetic[ICLR23] Uniformed Task
Ties-Merging[NeurIPS23] Uniformed Parameter
DARE[ICML24] 1/(1p)1/(1-p) Parameter
LoraHub[COLM24] Evolver Searched Task
AdaMerging[ICLR24] Unsupervised Optimized Layer
PCB-Merging[NeurIPS24] Balancing Matrix Parameter
NeuroMerging (Ours) L1-Norm Neuron
Table 2: Different merging scales and granularity levels.
@@
Model merging began with weight interpolation Wortsman et al. (2022) to combine the strengths of different models in weight space by balancing competition and cooperation within shared representation Du et al. (2024); Ilharco et al. (2023); Ortiz-Jimenez et al. (2023). In NLP, methods such as merging task-specific language models have been explored to build and update foundation models with multi-task capabilities Wan et al. (2024a); Akiba et al. (2025); Wan et al. (2024b). Similarly, in CV, approaches like merging Vision Transformers (ViTs) trained on different tasks or domains have been investigated to create unified models capable of handling diverse visual tasks Kim et al. (2021); Bao et al. (2022); Wang et al. (2024). In the multi-modal space, model merging has been applied to integrate models from different modalities, such as text and images, enhancing tasks like audio-visual question answering and image captioning Sung et al. (2023); Sundar et al. (2024); Dziadzio et al. (2025). These advancements underscore model merging as a promising avenue for future research.
Existing methods for model merging primarily operate at three granularities: model-level, layer-level, or parameter-level Ilharco et al. (2023); Yang et al. (2024b); Du et al. (2024), while overlooking the fundamental roles of neurons, their activation and connectivity Suhaimi et al. (2022); Stelzer et al. (2021), which underpins the learning process all neural networks from Perception Rosenblatt (1958) to LLMs Touvron et al. (2023). Table 2 provided a summary of existing methods across different granularities. In Figure 2 and Table 6 , we analyze on how changes in model weights along two complementary neuronal subspaces, leads to distinct impacts on task performance, by removing one and retaining the other. Notably, one subspace preserves most of the task-specific capabilities. This observation motivates us to explore model merging at the neuronal level, which could have important implications for mitigating task interference and could yield more robust merged models.
@@
Figure 3: Illustration of our proposed framework. (A) Our approach explicitly considers neuronal activation mechanisms through neuronal task vectors $\tau_{t}^{k}\in R^{d}$, defined as the difference between fine-tuned and pre-trained neurons for task $t$. (B) These vectors are decomposed into parallel and orthogonal subspaces relative to pre-trained neurons, corresponding to input sensitivity and task adaptability. SVD constructs a coordinate system in the previously unstructured orthogonal subspace. (C) Our NeuroMerging method merges models efficiently within these low-dimensional complementary subspaces, rather than in the original high-dimensional weight space.
@@ 2 Related Work
Multi-Task Learning (MTL) Fifty et al. (2021) leverages transferable knowledge to handle multiple related tasks simultaneously. Existing MTL approaches primarily rely on architectural design or optimization strategies. Architectural-based methods, such as Mixture of Experts (MoE) Shazeer et al. (2017), introduce specialized subnetworks that dynamically route inputs to task-specific experts, effectively reducing interference. However, these methods require modifying the pretrained model structure, increasing computational complexity, and limiting scalability Liu et al. (2019); Shen et al. (2024); Lu et al. (2024). Optimization-based approaches, on the other hand, focus on balancing task gradients or loss functions to mitigate task conflicts during training Bai et al. (2023); Cipolla et al. (2018). While these methods improve convergence, they still depend on task-specific training data, which may be impractical in real-world applications due to privacy concerns or data scarcity Liang et al. (2022). In contrast, model merging offers an alternative paradigm by integrating knowledge from multiple fine-tuned models into a single unified model without requiring additional training data or architectural modifications Wortsman et al. (2022); Ilharco et al. (2023). Notwithstanding the promising findings, a key challenge in model merging is task conflict Yadav et al. (2023); Du et al. (2024), where different tasks compete for model capacity, potentially leading to suboptimal performance.
@@ 3 Methodology
In this section, we first formalize the concept of neuronal task vector for model merging and then decompose neuronal task vectors into two complementary neuronal subspaces. Subsequently, we introduce our framework, NeuroMerging, shown in Algorithm 1 in Appendix A.4, which performs merging in the neuronal subspaces.
@@ 3.4 NeuroMerging
where $\psi_{\parallel}(\cdot)$ denotes the merging function for $\tau^{k}_{\parallel,t}\in\mathbb{R}$, which can be a commonly used weighted average, TIES’s disjoint merge Yadav et al. (2023), and others. $\psi_{\parallel}(\cdot)$ is also applied to non-neuronal parameters such as bias and pre-norm. For $\tau^{k}_{\perp,t}\in\mathbb{R}^{T}$, we first find dominant orthogonal subspaces within $\mathcal{O}$ using singular value decomposition (SVD) with rank being the number of tasks interact within the same neuron, and then project $\tau^{k}_{\perp,t}$ along each dimension of the SVD subspace, before applying $\psi_{\parallel}(\cdot)$ to them. Detailed discussion on SVD is provided in Appendix A.3.
$\lambda_{1}$ and $\lambda_{2}$ are scaling parameters. When validation data is available $\lambda_{1}$ and $\lambda_{2}$ could be tuned. However, when validation is unavailable, we propose setting $\lambda_{1}=0$ (as the corresponding subspace is observed to have little impact in Section 6.3) and estimate $\lambda_{2}$ based on the impact of top $r\%$ mentioned in Section 3.1. Specifically, $\lambda_{2}$ is estimated as $\lambda_{2}=\frac{1}{1-\sigma}$, where $\sigma=max(\sigma_{1},...,\sigma_{t},...,\sigma_{T})$ and $\sigma_{t}=\frac{\|\mathbf{\tau}_{t}^{masked}\|_{1}}{\|\mathbf{\tau}_{t}\|_{1}}$ is the ratio of the $L_{1}-Norm$ of the zeroed-out elements in the task vector $\mathbf{\tau}_{t}^{masked}$ to the $L_{1}-Norm$ of the original task vector $\mathbf{\tau}_{t}$. Detailed discussion on the parameters is provided in Section 6.3.
Subsequently, we reconstruct the task vector $\overline{\tau}$ with the all the merged neuronal task vectors $\overline{\tau}_{k}$. Finally, we obtained the final merged model $\overline{\theta}=\theta_{0}+\overline{\tau}_{rescaled}$. The complete procedure is presented in Algorithm 1, with additional discussion on the motivation provided in Appendix B.
@@ 4 Experimental Setup
Baseline Methods. Our baselines comprise two main categories: (1) non-model merging approaches, which include individually fine-tuned models and a multitask model trained jointly on the combined dataset serving as our theoretical upper bound, and (2) various advanced model merging techniques, including Simple Averaging (Wortsman et al., 2022), Fisher Merging (Matena and Raffel, 2022), RegMean (Jin et al., 2023), Task Arithmetic (Ilharco et al., 2023), TIES-Merging (Yadav et al., 2023), AdaMerging (Yang et al., 2024b), PCB-Merging (Du et al., 2024), NPS (Du et al., 2025), and our proposed NeuroMerging method. Notably, we used task-wise AdaMerging for fair comparison with existing baselines and our method. See Section 6.5 for further discussion and analysis. We reported average accuracy across all tasks’ test sets as our primary evaluation metric.
Validation Set Availability. Previous works exhibited varying dependencies on a validation set. Fisher Merging inherently required a validation set to compute the Fisher matrix. Other approaches may optionally utilize validation data for hyperparameter tuning, while RegMean leverages training data to compute and store inner product matrices for model merging. However, since these matrices matched the dimensions of the original model, they introduced substantial storage and computational overhead, limiting scalability to larger models and more extensive merging tasks.
@@
Hyperparameters. In the absence of an additional validation set, we set $\lambda=1$ as the default value for all task-vector-based methods. For TIES-Merging and PCB-Merging, which required a masking ratio, we followed the settings of Yadav et al. (2023) and Du et al. (2024), applying $r=0.2$ as the default value across all experiments. For NeuroMerging, we set a default masking ratio of $r=0.15$, with $\lambda_{1}$ fixed at 0, while $\lambda_{2}$ was automatically adjusted according to the methodology described in Section 3.4.
When validation is allowed, we configure the non-diagonal multiplier $\alpha$ in RegMean to $0.9$, except for the T5-base model, where it is set to $0.1$. For Task Arithmetic, we performed a grid search over $\lambda$ ranging from $0.2$ to $1.5$ with a step size of $0.1$. For TIES-Merging, PCB-Merging, and NeuroMerging, we searched for the optimal masking ratio r in the range $[0.05,0.2]$ with a step size of $0.05$, and $\lambda$ ($\lambda_{2}$ for NeuroMerging) from $0.8$ to $5.0$ with a step size of $0.1$.
@@ 5 Results
5.1 Merging NLP Models
Following the experimental settings from Yadav et al. (2023), we used the T5-base and T5-large models (Raffel et al., 2020), which were encoder-decoder transformers (Vaswani et al., 2017) pretrained via masked language modeling on a large text corpus, and fine-tuned them independently on seven tasks: Khot et al.’s (2020) QASC, Yang et al.’s (2015) WikiQA, and Tafjord et al.’s (2019) QuaRTz for Question Answering; Zhang et al.’s (2019) PAWS for Paraphrase Identification; Sharma et al.’s (2018) Story Cloze for Sentence Completion; and Sakaguchi et al.’s (2020) Winogrande together with Levesque et al.’s (2012) WSC for Coreference Resolution. Tables 3 and 13 demonstrated that our approach achieved superior performance over state-of-the-art methods, achieving improvements of $0.6\%$ and $1.2\%$ for T5-base and T5-large, respectively. Moreover, NeuroMerging without validation showed even more substantial gains, surpassing previous methods by $1.4\%$ for T5-base and $0.7\%$ for T5-large. For comprehensive results across all tasks and model variants, see Appendix C.1 Tables 3 and 13, with detailed error analysis provided in Appendix C.4.
5.2 Out-of-Domain Generalization
Building upon the experimental setup of Yadav et al. (2023), we also investigated how merged models between tasks enhance generalization in different domains. Following the approach used in prior NLP models, we merged models on seven in-domain datasets and evaluated their performance on six held-out datasets from the T0 mixture (Sanh et al., 2022) to assess cross-task generalization. These datasets encompass diverse tasks, covering three Question Answering datasets: Huang et al.’s (2019) Cosmos QA, Sap et al.’s (2019) Social IQA, and Rogers et al.’s (2020) QuAIL; one Word Sense Disambiguation dataset: Pilehvar and Camacho-Collados’s (2019) WiC; and two Sentence Completion datasets: Gordon et al.’s (2012) COPA and Zellers et al.’s (2019) H-SWAG. As shown in Figure 4, NeuroMerging outperformed the strongest baseline by $0.7\%$ and $0.5\%$ for T5-base and T5-large models, respectively, showcasing enhanced out-of-domain generalization capabilities. For more comprehensive results, please refer to Appendix C.1 Tables 16 and 17, with detailed error analysis provided in Appendix C.4.
@@ 5.3 Merging LLMs
We followed the experimental setup of Du et al. (2024) and extended our NeuroMerging to larger LLMs. Specifically, we merged three domain-specialized Llama-2-7b models (Touvron et al., 2023), and each was fine-tuned for distinct capabilities: Chinese language understanding, mathematical reasoning (Yu et al., 2024b), and code generation (Roziere et al., 2023). To rigorously evaluate the performance of each specialized model, we employed established benchmarks tailored to their respective domains: Li et al.’s (2024) CMMLU for assessing Chinese language proficiency, Cobbe et al.’s (2021) GSM8K for mathematical capabilities, and Chen et al.’s (2021) HumanEval for code generation competency. As demonstrated in Table 4, our approach exhibited substantial performance improvements, surpassing the strongest baseline by $0.9\%$. Notably, our method even outperformed the PCB-Merging utilizing Evolution Strategies (ES) optimization algorithm by $0.7\%$, underscoring the effectiveness of our proposed methodology. Appendix C.4 provides the error analysis.